OpenStudy (anonymous):

examine the following sets for linear independence: u1=eix, u2=e-ix, u3=sinx

OpenStudy (nowhereman):

sorry, but those are no sets.

OpenStudy (anonymous):

\[u_{1}=e ^{ix}, u _{2}=e ^{-ix}, u _{3}=sinx\]

OpenStudy (nowhereman):

still those are only terms with the free variable x

OpenStudy (nowhereman):

do you mean \[\{e^{ix}, e^{-ix}, \sin{x}\}\] for an arbitrary x or do you mean \[\{e^{ix} | x \in ℂ \}\] and so on?

OpenStudy (anonymous):

an arbitrary

OpenStudy (anonymous):

\[c _{1}u _{1}+c _{2}u _{2}+c _{3}u _{3}\] if linear combination is zero c1, c2, c3=0 then its called linear independence.

OpenStudy (nowhereman):

they dependent: \[\sin x = \frac{e^{ix} - e^{x}}{2i}\]

OpenStudy (anonymous):

could you explain how this combination come?

OpenStudy (nowhereman):

It arises naturally if you define sinus and the exponential function by their power series. But also if you take geometric intuition and define \[e^{a+bi} = e^a(\cos b + i \sin b)\]